Cassini Significant Events 11/03/10 - 11/09/10
Cassini Significant Events 11/03/10 - 11/09/10
Nov. 11, 2010
Wednesday, Nov. 3 (DOY 307)
On Tuesday, DOY 306, the Spacecraft Operations team was performing a real time command procedure to load the backup AACS Flight Computer-A (AFC-A) with the new AACS A8.8.0 flight software from the Solid State Recorder (SSR) when a sudden loss of downlink occurred, halting the procedure. The DSN station at Canberra, Australia, DSS-45, provided support during the loss of downlink and confirmed execution of High Gain Antenna (HGA) safing. It was later determined that a command had been corrupted by a flipped bit and caused a non-maskable interrupt error in the prime Command and Data Subsystem A (CDS-A). CDS-A performed a reset which correctly resulted in CDS-B becoming prime. The spacecraft safing algorithm was run, which terminated all active programs and shut down all nonessential power loads, including all science instruments, and resulting in deactivation of the executing S64 background sequence. The spacecraft then turned to Sun-point and switched to the low gain antenna. This caused DSS-45 to lose lock on the downlink signal at 2010-306T23:56 UTC / Tuesday, Nov. 2 at 4:56 pm PST. An hour later, the HGA Swap response executed and communication was restored through the high gain antenna and DSS-45 was able to acquire the 1896 bps engineering telemetry downlink.
Recovery efforts continued today, with 12 command files being sent to the spacecraft. These files set the uplink rate back to 500 bps, turned on ranging, powered on the Visual and Infrared Mapping Spectrometer (VIMS) instrument to get it to a safe thermal state, turned on the prime pressurant control assembly line heater, enabled Autonomous Thermal Control #6, enabled the HGA Swap monitor, cleared system fault protection statistics, and read out various AACS and CDS memories for diagnostic purposes.
Thursday, Nov. 4 (DOY 308)
Recovery efforts continued today with six command files being sent to the spacecraft. These files were sent to play back engineering data around the time of the safing incident, read out more CDS memory, clear AACS high watermarks, and open propulsion latch valve 40, which had been closed by the safing event. The commands also cleared system fault protection and CDS error logs, and put on board a mini-sequence to control the telemetry modes until the S65 background sequence goes active on Nov. 24.
Cassini personnel continue to support the office moves within the Project area. So far, 26 out of 60 office moves have been completed. This effort is scheduled to continue through next month, when it's projected that all 60 moves will be complete. This is an ongoing task driven by the Project's staff reduction as part of its transition to the new, descoped Solstice mission and the accommodation of other Projects moving into the operations building.
Friday, Nov. 5 (DOY 309)
The working group for the Distributed Object Manager (DOM) transition to the Network File System (NFS) has successfully set up a test DOM NFS server inside the firewall, ready for testing except for those outside of the firewall. Currently, DOM is on the legacy Andrew File System (AFS), a distributed network file system which is no longer supported by the Laboratory. This effort to move the DOM from AFS to NFS is still ongoing and slated for completion early in 2011.
Monday, Nov. 8 (DOY 312)
Orbit Trim Maneuver (OTM) #265 was performed today. This was a Reaction Control Subsystem (RCS) burn which began at 3:15 AM PST. Telemetry immediately after the maneuver showed a burn duration of 157.5 seconds, giving a total delta-V of 176.92 mm/sec, as planned. This was executed using the all RCS maneuver block which means the roll and yaw turns to and from the maneuver attitude were done under RCS control instead of the normal reaction wheel control as a consequence of the earlier spacecraft safing event. This was the first OTM to be done using a pmxfile interface in place of the Maneuver Operations Program Set/Maneuver Automation Software (MOPS/MAS) listener interface. The switch to pmxfiles was done in preparation for the Navigation Team's switch to the MONTE software set next year. All subsystems reported nominal performance after the maneuver.
Recovery efforts today included uplinking a file with AACS vectors and clear high water mark commands and a file to deregister seven programs from the new prime CDS-B.
A status update titled "Cassini to Resume Normal Operations" was released today and posted on the Cassini web site. Engineers at JPL expect the Cassini spacecraft to resume normal operations on Nov. 24. They have traced the steps taken by an onboard computer when Cassini put itself in precautionary 'safe mode' last week. Engineers determined that the spacecraft went into safe mode because of a flip of a bit in the command and data system computer. For more information on this subject link to:
Tuesday, Nov. 9 (DOY 313)
Cassini supported telemetry testing of the DSN's new Data Capture and Delivery (DCD) software by participating in DSN-DCD Project Interface Tests (PIT) in mid-September and late October. Review of DCD PIT results revealed that average latency for Cassini increases for DCD, and that some gaps in data are difficult to determine because of different hardware paths. Further analysis to explain latency and data issues is still ongoing. DCD installation at the DSN complexes is scheduled to start today with the upgrade at the Madrid, Spain complex, to be completed by Nov. 10. The Goldstone complex is tentatively scheduled for the DCD upgrade on Nov. 29, followed by the Canberra complex on Dec. 14.As part of the Mission Control, Data Management and Accountability and Spacecraft Analysis (MDAS) 5.0.5 delivery, a fix to correct the lack of notifications when valid telemetry red alarms occur was included. Cassini supported testing of this new AAN v32/Solaris-10 version, which was tested successfully in operations last Wednesday.